# DeepCloth-ROBP&P: Towards a Robust Robot Deployment for Quasi-Static Pick-and-Place Cloth-Shaping Neural Controllers

### v1

@@ IV-A Generasibility of Towel-Sim2Real
To demonstrate generasability of our deployment protocol Towel-Sim2Real, we conducted experiments with 4 different neural controllers and 6 different fabrics on 4 different shaping tasks. Each physical flattening trial concluded within a maximum of 20 steps normalised coverage exceeding 95% defined as success. We also examine the neural controllers on all-corner inward folding (4 steps), corners-edge inward folding (4 steps) and diagonal-cross foldings (2 steps) with extra 2 steps to allow controllers to recover from mistakes. Intersection-over-Union metric between current and goal masks is used to evaluate step-wise performance, with 85% considered a success. For each shaping task, we also verify success through human judgement. As a control, we also provide a human baselines with a robot executing actions entered via a pick-and-place human interface.
@@
We train PlaNet-ClothPick, JA-TN and diffusion policies in the newly proposed Real2Sim Towels benchmark environment — PlaNet-ClothPick with 20 thousands transitional steps (instead of its original 1 million) using its specially engineered collection policy [5] and the later two with 2000 trajectories created by our updated oracle. Evaluation for flattening controllers in physical trials typically employs step-wise normalised coverage, normalised improvement (requiring real-time cloth segmentation), as well as operational efficiency [1, 3]. Table I shows that all neural policies can transfer successfully to real world scenarios for Fabrics C and D, but all struggle with flattening Fabric A. Although JA-TN performs worse than diffusion policy in simulation, JA-TN does slightly better in real-world in general. Compared to human performance, all methods still needs improvement especially with really soft fabrics like Fabric A.
@@
On the other hand, assessing folding tasks is more complex due to task variety and the self-occlusive nature of goal states, often necessitating human judgement. Intersection-over-Union (IoU) serves as a reasonable proxy metric for step-wise folding performance in real world [7, 10, 4]. The initial state of the fabric are required to be wholly inside the camera view. We only test on the fabrics with low stiffness, such as Fabric A and B, so they will not flip back to their last state.
@@
We retrain JA-TN and diffusion policies with 1000 oracle-based demonstrations [11] in Real2Sim Towels. We also adopt the off-the-shelf Foldsformer model with its original deployment strategy as the baseline for folding comparison. Table II demonstrates that both JA-TN and Diffusion Policy can deploy successfully in reality with our Towel-Sim2Real protocol. Diffusion approaches human performance whilst JA-TN and Foldsformer perform similarly and slightly worse. Despite performing well in all-corner-inward and diagonal-cross foldings, all methods mainly struggle in more complex corners-edge inward folding.
@@
Tables I and II empirically demonstrate that our Towel-Sim2Real training and deployment protocol applies across various neural controller and towel-shaping tasks. It also shows room for improvement on flattening soft fabrics (Fabric A) and performing corners-edge inward folding tasks.
@@
IV-B Robustness of Towel-Sim2Real
Towel-Sim2Real includes three key components: (1) the Real2Sim Towels environment along with its oracle flattening policy R2S-S, (2) a new cloth segmentation heuristics for masking vision inputs and reliable grasping for fabrics, and (3) the image processing for training and deployment proposed in JA-TN [11]. Figure 2 illustrates the reliability of our cloth segmentation method. To assess the significance of the other two components, we compare 4 neural controllers with depth-only input, and we choose Fabric C and D that all policies did well in Table I. We train JA-TN, OTS in Rainbow Rectangular Fabrics and JA-TN, R2S-S in Real2Sim Towels both using JA-TN’s vision processing which also applied to the Diffusion policy. We train another Diffusion, R&S Height with masked-depth input without special processing deployed by adding the difference of simulated and real camera.
Figure 3 shows that the changes in the R2S-S oracle policy represent a substantial leap towards more robust and practical cloth manipulation in real-world robotic applications. It indicates that the diffusion policy trained and deployed with Towel-Sim2Real shows dominant performance on real fabrics, whilst JA-TN, R2S-S and Diffusion, R&S Height only reaches similar performance in simulation. The policies trained with R2S-S demonstration using JA-TN’s vision processing strategy shows better performance that other policies that uses either simple vision processing strategy or trained with OTS policy.
IV-C Grasping with Tweezers-extended Gripper
We evaluate the tweezers-extended gripper against the default Franka gripper aiming to minimise misgrasping, multilayer grasping, cloth dropping and cloth attachment. We encompass a diverse range of cloth conditions including three cloth types (silk, cotton, and polyester), various cloth states (Flat, 2-folds, 3-folds), and multiple grasping scenarios (4 corners, 4 edges, and 2 random cloth points). Table III indicates that the tweezers perform better compared to the Franka gripper across all pick types, particularly excelling in complex graspings on four edges and corners with over 90% success rate. Misgrasp rates for the tweezers remain consistently below 10%, while the Franka Gripper shows higher rates between 16-22%. Both grippers perform similarly for multilayer handling, but the tweezers completely avoid cloth drops and attachment issues, which occur more frequently with the default gripper. Furthermore, the tweezers maintain consistent performance across all cloth types, with particular excellence in handling delicate fabrics like silk.
V Discussion
Many neural controllers in towel shaping only work in specific cloth characteristics regarding particular colour, thickness, and initial cloth positions in camera view [1, 4, 7]; their generalisability to other fabric contexts and robot platform is uncertain. With our protocol DeepCloth-ROBP&P, we deployed four neural policies, including a diffusion policy, in at least two robotic platforms, including Franka Emika Panda with an ”eye-in-hand” setting and UR3e with a ”hand-to-eye” setting.
@@ V Discussion
Soft fabrics with high internal friction pose challenges for flattening methods due to the difficulty in separating different parts of the fabric. Conversely, small and stiff fabrics are problematic for folding with a single end-effector as they tend to flip back easily. For all folding controllers, the initial state of the folding tasks should be completely inside the field of vision. Our current robotic system faces several limitations: (1) it struggles to automatically detect success states accurately; (2) it occasionally produces helical and spiral path trajectories; and (3) it still experiences minor misgrasping issues. Fine-tuning camera extrinsic for precise pixel-to-base conversion also remains tedious and time-consuming. Identifying an appropriate workspace for smooth and stable pick-and-place operations is also tricky, primarily due to the robot’s limited reach and the tendency of planned trajectories to produce erratic joint movements when approaching the robot base. Last but not the least, we find that the early checkpoints of the trained networks deploys better for most models, but the surprising diffusion policy does not overfit to the artefact of the simulation under Towel-Sim2Real.
@@
We plan to apply the proposed method to more difficult tasks, such as flattening and folding different garment types. We want to further investigate the capacity of diffusion policy through making folding agnostic to the cloth’s initial state and achieve folding-from-crumpled in the real world. Integrating depth estimation along with a tactile sensor may help to improve the dexterity of the grasping and substantially reduce the effect of grasping multiple layers of the articles, which may be vital for the successful manipulation of complex garment manipulation.
VI Conclusion
We introduce a set of robust robotic strategies, collectively known as DeepCloth-ROBP&P, that enable effective real-world deployment of pick-and-place cloth-shaping neural controllers. We comprehensively evaluate four different neural controllers’ real-world performance on shaping towels with varying size, colour, shape, material and thickness, including a modified visuomotor diffusion policy. Our method also demonstrates the versatility of the proposed protocol across different platforms. Ultimately, this research contributes valuable insights and practical solutions to robotic cloth manipulation, paving the way for more advanced and reliable systems in real-world cloth manipulation.
VII Acknowledgement

### v2

@@ V-A Generalisability of DRAPER
To demonstrate generalisability of our deployment framework DRAPER, we conduct experiments with 4 different neural controllers and 6 different fabrics on 4 different shaping tasks. We also examine the neural controllers on all-corner inward folding (4 steps), corners-edge inward folding (4 steps) and diagonal-cross foldings (2 steps) with extra 2 steps to allow controllers to recover from mistakes. As a control, we also provide a human baselines with a robot executing actions entered via a pick-and-place human interface.
We train PlaNet-ClothPick, JA-TN and diffusion policies (50, 80 and 100 thousand updates respectively) with RGB as input and DRAPER’s training protocol. Note that PlaNet-ClothPick is trained on a dataset with 20 thousands transitional steps (instead of its original 1 million) using its specially engineered collection policy [5] and the latter two policies with 2000 successful trajectories created by our new real-world adapted Oracle policy, RealAdapt-OTS, so that all policies will have a similar amount of training data.
@@
Evaluation for flattening controllers in physical trials typically employs step-wise normalised coverage, normalised improvement (requiring real-time cloth segmentation), as well as operational efficiency [1, 3]. Table I shows that all neural policies can transfer successfully to real-world scenarios for Fabrics C and D, but all struggle with flattening Fabric A. JA-TN performs the best in real-world in general. Compared to human performance, all methods still need improvement especially with soft fabrics like Fabric A.
@@
On the other hand, assessing folding tasks is more complex due to task variety and the self-occlusive nature of goal states, often necessitating human judgement. IoU serves as a reasonable proxy metric for step-wise folding performance in real world [7, 10, 4]. The initial state of the fabric are required to be wholly inside the camera view. We only test on the fabrics with low stiffness, such as Fabric A and B, so they will not flip back to their last state. We train JA-TN and diffusion policies with 1000 oracle-based demonstrations [8] in in our newly proposed simulation environment RealAdapt Towels. We also adopt the off-the-shelf Foldsformer model with its original deployment strategy as the baseline for folding comparison.
@@
Table II demonstrates that both JA-TN and diffusion folding policy can be deployed successfully in the real world with our DRAPER framework. Diffusion policy shows near human performance whilst JA-TN and Foldsformer demonstrate similar performance to each other yet slightly inferior to the diffusion policy. Despite performing well in all-corner-inward and diagonal-cross foldings, all methods mainly struggle with more complex corners-edge inward folding.
@@
Tables I and II empirically demonstrate that our DRAPER framework results in robust deployment of various neural controller to achieve towel-shaping tasks in the real world. It also shows room for improvement on flattening soft fabrics (Fabric A) and performing corners-edge inward folding tasks.
@@
V-B Ablation Study of DRAPER
In this ablation study, we aim to study the effect of different components of the DRAPER framework. We conduct this investigation in both simulation and real-robot setups. Unlike the previous section, we train all diffusion policies with 50,000 updates to clearly demonstrate the effect of the framework.
For the simulation evaluation, as shown in Figure 3, we create a new testing environment, RealAdapt Towels (No Misgrasp), where we disable the misgrasping scenarios, so that the actual ability of these policies can be demonstrated accurately. Additionally, the real-world experiments, as shown in Table III, indicate that all components of DRAPER are critical for reliable deployment of the neural controllers. Also, training purely on depth inputs and doubling the training steps of RGB-D diffusion policy can help improve performance across different simulation settings.
V-C Grasping with Tweezers-extended Gripper
We evaluate the tweezers-extended gripper against the default Franka gripper aiming to minimise misgrasping, multilayer grasping, cloth dropping and cloth attachment. We encompass a diverse range of cloth conditions including three cloth types (silk, cotton, and polyester), various cloth states (Flat, 2-folds, 3-folds), and multiple grasping scenarios (4 corners, 4 edges, and 2 random cloth points). Table IV indicates that the tweezers perform better compared to the Franka gripper across all pick types, particularly excelling in complex graspings on four edges and corners with over 90% success rate. Misgrasp rates for the tweezers remain consistently below 10%, while the Franka Gripper shows higher rates between 16-22%. Both grippers perform similarly for multilayer handling, but the tweezers completely avoid cloth drops and attachment issues, which occur more frequently with the default gripper. Furthermore, the tweezers maintain consistent performance across all cloth types, performing particularly well with delicate fabrics like silk.
VI Discussion
Many neural controllers in towel shaping have only been tested with cloths of particular colour, thickness, and initial positions in camera view [10, 8, 5]; their generalisability to other fabric contexts and robot platform is uncertain. Our DRAPER framework allows us to deploy four neural policies, including a diffusion policy, on two different robotic platforms: Franka Emika Panda with an ”eye-in-hand” setting and UR3e with a ”hand-to-eye” setting.
Through our experiments, we find that soft fabrics with high internal friction pose challenges for flattening methods due to the difficulty in separating different parts of the fabric. Conversely, small and stiff fabrics are problematic for folding with a single end-effector as they tend to flip back easily.
@@ VI Discussion
Although we successfully deploy numerous simulation-trained neural controllers in real-world setups, our current robotic systems, i.e., the tested neural controllers with DRAPER real-world deployment, faces several limitations: (1) it struggles to automatically detect success states accurately; (2) it occasionally produces helical and spiral path trajectories; (3) it still experiences minor misgrasping issues; and, (4) for all folding controllers, the initial state of the folding tasks required to be completely inside the field of vision.
@@
We plan to apply the proposed method to more difficult tasks, such as flattening and folding different garment types, such as T-shirts and trousers. We want to further investigate the capacity of diffusion policy through making folding agnostic to the cloth’s initial state and achieve folding-from-crumpled in the real world. Integrating depth estimation along with a tactile sensor may help to improve the dexterity of the grasping and substantially reduce the effect of grasping multiple layers of the articles, which may be vital for the successful manipulation of complex garment manipulation.
VII Conclusion
We introduce a set of robust robotic deployment techniques, collectively known as DRAPER deployment framework, that enables effective real-world deployment of pick-and-place cloth-shaping neural controllers. We perform the first comprehensive comparison of four SotA neural controllers in real-world settings, evaluating shaping of towels with varying size, colour, shape, material and thickness. This includes the first application of a visuomotor diffusion policy on this task. Our method also demonstrates the versatility of the proposed protocol across different robotic platforms. Ultimately, this research contributes valuable insights and practical solutions to robotic cloth manipulation, paving the way for more advanced and reliable systems in real-world cloth manipulation.
VIII Acknowledgement